40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 14:42:34 2011, steve, Update, SUS, oplev gains today 
    Reply  Wed Jul 20 19:24:12 2011, rana, Update, SUS, oplev gains today Untitled.pnga.png
       Reply  Wed Jul 20 19:48:03 2011, Jamie, Update, SUS, Re: oplev gains today 
       Reply  Thu Jul 21 09:04:59 2011, valera, steve, Update, SUS, oplev gains were not optimized 
       Reply  Mon Jul 25 00:35:44 2011, rana, Update, SUS, something wrong with ETMY LR sensor a.pngbad.png
          Reply  Tue Jul 26 11:38:45 2011, steve, Update, SUS, ETMY bias checked 
    Reply  Thu Jul 21 16:06:35 2011, steve, Update, SUS, oplev gains today checked by conlog 
Message ID: 5001     Entry time: Wed Jul 20 14:42:34 2011     Reply to this: 5004   5014
Author: steve 
Type: Update 
Category: SUS 
Subject: oplev gains today 

C1:SUS-ETMX_OLPIT_GAIN set to 1.0     OLYAW 1.0

                ETMY                                     -0.2                   -0.2

                ITMY                                        2.0                    -2.0

                ITMX                                        0.5                     0.5

                  BS                                          0.4                    -0.4

                PRM                                         0.5                    -0.7

                SRM                                         1.0                     1.0

Earlier today Rana and I made power spectra of ETMY_OPLEV_ERROR signals with servo on and off.

It was indicating that the servo is not doing anything. These gain values were not set since IFO rebuilt.

Valera's entries were searched also. He did not do such thing. Rana may know where it is in the elog if it happened.

ELOG V3.1.3-